# HiLo: A Learning Framework for Generalized Category Discovery Robust to Domain Shifts

### v1

@@ 4.1 Experimental setup
Datasets. To validate the effectiveness of our method, we perform various experiments on the largest public datasets with domain shifts, DomainNet [31], containing about 0.6 million images with 345 categories distributed among six domains. Moreover, based on the Semantic Shift Benchmark (SSB) [29] (including CUB [76], Stanford Cars [77], and FGVC-Aircraft [78]), we construct a new corrupted dataset called SSB-C (i.e., CUB-C, Scars-C, and FGVC-C) following [30]. We exclude unrealistic corruptions and corruptions that may lead to domain leakage to ensure that the model does not see any of the domains in SSB-C during training (see Appendix A for details). Overall, we introduce 9 types of corruption and 5 levels of corruption severity for each type, resulting in a dataset 45$\times$ larger than SSB.
@@
For the semantics axis, on both DomainNet and SSB-C, following [8], we sample a subset of all classes as the old classes and use 50% of the images from these labelled classes to construct $\mathcal{D}_{\Omega^{a}}^{l}$. The remaining images with both old classes and new classes are treated as the unlabelled data $\mathcal{D}_{\Omega^{a}}^{u}$. For the domain axis, on DomainNet, we select images from the ‘Real’ domain as $\mathcal{D}_{\Omega^{a}}$ and pick one of the remaining domains as $\mathcal{D}_{\Omega^{b}}$ in turn (or include all the remaining domains as $\mathcal{D}_{\Omega^{b}}$). While on SSB-C, we use each dataset in SSB as $\mathcal{D}_{\Omega^{a}}$ and its corresponding corrupted dataset in SSB-C as $\mathcal{D}_{\Omega^{b}}$. Statistics of the datasets are shown in Table 1.
@@
Evaluation protocol. For DomainNet, $\Omega^{a}=\left\{\omega_{1}\right\}$ and $\Omega^{b}=\left\{\omega_{2}\right\}$, where $\omega_{i}$ stands for different domains. We also experiment with the case where $\Omega^{b}=\left\{\omega_{2},\cdots,\omega_{6}\right\}$. We train the models on $\mathcal{D}_{\Omega^{a}}$ (i.e., $\mathcal{D}_{\Omega^{a}}^{l}\cup\mathcal{D}_{\Omega^{a}}^{u}$) and $\mathcal{D}_{\Omega^{b}}^{u}$ of all classes without annotations. For SSB-C, $\Omega^{a}=\left\{\omega_{1}\right\}$ and $\Omega^{b}=\left\{\omega_{2},\cdots,\omega_{10}\right\}$ since we have nine types of corruptions. During evaluation, we compare the ground-truth labels $y_{i}$ with the predicted labels $\hat{y}_{i}$ and measure the clustering accuracy by $ACC=\frac{1}{|D^{u}|}\sum_{i=1}^{|D^{u}|}\mathds{1}(y_{i}=\phi(\hat{y}_{i}))$, where $\phi$ is the optimal permutation that matches the predicted cluster assignments to the ground-truth labels. We report the ACC values for ‘All’ classes (i.e., instances from $\mathcal{Y}$), the ‘Old’ classes subset (i.e., instances from $\mathcal{Y}^{l}$), and ‘New’ classes subset (i.e., instances from $\mathcal{Y}^{u}$) for $\mathcal{D}_{\Omega^{a}}^{u}$ and $\mathcal{D}_{\Omega^{b}}^{u}$ separately.
@@
Implementation details. Following the common practice in GCD, we use the DINO [36] pre-trained ViT-B/16 as the feature backbone and the number of categories is known as in [5] for all methods for fair comparison. When the category number is unknown, one can employ existing methods (e.g., [1, 8, 38, 42]) to estimate it and substitute it into the category discovery methods (see Table 24). The 768-dimensional embedding vector corresponding to the CLS token is used as the image feature. For the feature backbone, we only fine-tune the last Transformer layer. We train each dataset for $T=200$ epochs using a batch size of 256. We follow the protocol in  [8, 5] to select the optimal hyperparameters for our method and all baselines, based on the ‘All’ accuracy on the validation split of $\mathcal{D}_{\omega_{1}}$. The initial learning rate for our approach is 0.1 for CUB and 0.05 for other datasets, and the rate is decayed using a cosine schedule. $t^{\prime}$ is set to the $80$-th epoch. $r_{0}$ is assigned as $|\mathcal{D}^{l}|/|\hat{\Omega}^{b}|$ for DomainNet and 0 for SSB-C. $r^{\prime}$ is set to 1 for DomainNet and 0.05 for SSB-C. $\epsilon$ is set to 0.1. Following [8, 5], we set $\lambda=0.35$. See Appendices O and P for choices of hyperparameters for HiLo components and learning rates for all methods.
@@ 4.2 Main comparison
We compare our method with ORCA [37], GCD [8] and SimGCD [5] in generalized category discovery, along with two strong baselines RankStats+ [3] and UNO+ [4] adapted from novel category discovery, on DomainNet (Table 2) and SSB-C (Table 3), respectively. Additionally, we provide results by incorporating UDA techniques in Appendix E and the strong CLIP model in Appendix H.
In Table 2, we present results on DomainNet considering one domain as $\Omega^{b}$ each time. Our method consistently outperforms other methods for ‘All’ classes (even better for ‘New’ classes) in both domain $\Omega^{a}$ and $\Omega^{b}$ by a large margin. For example, for the ‘Real’ and ‘Painting’ pair, it outperforms the GCD SoTA method, SimGCD, by nearly 5% and 19% in proportional terms, which is remarkable considering the gap between different methods. RankStat+ performs well on ‘Old’ categories in the unseen domain $\Omega^{b}$. In Appendix B, we present results on DomainNet considering all domains except ‘Real’ as $\Omega^{b}$. It can be seen, in such a challenging mixed domain scenario, our method still substantially outperforms other methods. A breakdown evaluation of each domain shift for both datasets can be found in Appendices I and J.
@@
Table 3: Evaluation on SSB-C datasets. We report results of baselines in the seen domain (i.e., Original) and the overall performance of different corruptions (i.e., Corrupted). On ‘Corrupted’, our model provides between 20% and 80% relative gains over SimGCD [5].
@@
Comparing the results with the single domain results in [8, 5], we find that including corrupted data during training impairs the performance on the original domain. SSB-C is 45$\times$ larger than SSB, posing a significant challenge and resulting in unsatisfactory performance for existing methods. However, our method, HiLo, continues to demonstrate promising results, further validating its effectiveness.
@@
Reference SimGCD [5] 61.3 77.8 52.9 34.5 35.6 33.5
(1) SimGCD + PatchMix in [17] 62.5 76.3 54.2 34.8 36.0 33.8
@@
(5) Deep features only 28.2 40.3 22.7 13.6 20.0 11.0
(6) Shallow features only 10.1 18.1 6.4 5.7 9.2 5.7
@@ 4.3 Analysis
Effectiveness of different components. We validate the effectiveness of different components and design choices for our method in Table 4. As our method is built upon SimGCD, the effectiveness of each component can be observed by comparing its performance with that of SimGCD. We combine SimGCD with the original PatchMix in [17] (row 4) as a strong baseline for our task since these are SoTA methods for GCD and UDA respectively. Rows 4-4 indicate our main conceptual methodological contributions. As can be seen, simply combining SimGCD with the original PatchMix developed for UDA leads to a relatively small influence on the results. The original PatchMix focuses mainly on bridging the domain gap of labelled classes through a semi-supervised loss, which limits its capability on the unseen classes from new domains. After sutbly adapting PatchMix into contrastive learning for GCD (row 4), the unlabelled data containing both domain shifts and semantic shifts can be properly utilized for training, leading to an obvious performance boost on $\Omega^{b}$. Furthermore, when we disentangle semantic features from domain features (row 4), the model significantly improves performance on both $\Omega^{a}$ and $\Omega^{b}$, demonstrating dissociation of spurious correlations. Curriculum sampling further enhances performance on $\Omega^{b}$ (row 4).
@@
4.4 Qualitative results
Here, we provide qualitative results on DomainNet and CUB-C. The attention map offers valuable insights into the focus of Transformer-based models on the input. We obtain the attention maps for the CLS token from multiple attention heads in the final layer of the ViT backbone, highlighting the top 10% most attended patches in Figure 4. We observe that, compared with the baseline, HiLo is much more effective in focusing on the foreground object even in the presence of significant domain shifts (e.g., painting style, foggy weather). This demonstrates that HiLo is robust to domain shifts and remains unaffected by potential spurious correlations between semantic features and low-level statistics.
@@ Acknowledgements
This work is supported by Hong Kong Research Grant Council - Early Career Scheme (Grant No. 27208022), National Natural Science Foundation of China (Grant No. 62306251), and HKU Seed Fund for Basic Research.
@@ Appendix A SSB-C benchmarks
As demonstrated in Section 4.1 in the main paper, we construct the SSB-C benchmark to evaluate the robustness of algorithms to diverse corruptions applied to validation images of the original SSB benchmark (including CUB [76], Stanford Cars [79] and FGVC-Aircraft [78]), adopting the corruptions following [30]. We introduce 9 types of corruption (see Figure 5) in total. Each type of corruption has 5 severity levels. Therefore, SSB-C is 45$\times$ larger than the original SSB.
@@
Here is the list of the 9 types of corruption we applied following [30]:
@@
Though synthetic, SSB-C incorporates extra challenges and unique values over existing datasets like DomainNet [31]. Particularly, SSB-C includes (1) fine-grained recognition challenges under domain shifts and (2) more types of domain shifts that are not covered in DomainNet.
@@ Appendix D Theoretical analysis
Recall that $\mathcal{D}$ is an open-world dataset consisting of a labelled set $\mathcal{D}^{l}=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{N_{l}}\subset\mathcal{X}^{l}\times\mathcal{Y}^{l}$ and an unlabelled set $\mathcal{D}^{u}=\{\boldsymbol{x}_{i}\}_{i=1}^{N_{u}}\subset\mathcal{X}^{u}$. We also define the mapping function $g$ parametrized by a deep neural network as one of the hypotheses from $\mathbb{G}$.
@@
When the source domain does not adequately cover the target domain, the target risk of a learned hypothesis cannot be consistently estimated without additional assumptions. Nonetheless, an upper bound on the target risk can be estimated and then minimized. [80] introduce the $\mathcal{A}$-distance (also known as $\mathcal{H}$-divergence) to assess the worst-case loss when extrapolating between domains for hypothesis classes. The $\mathcal{A}$-distance between any two distributions $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ is defined as

### v2

@@ 4.1 Experimental setup
Datasets. To validate the effectiveness of our method, we perform various experiments on the largest public datasets with domain shifts, DomainNet Peng et al. (2019a), containing about 0.6 million images with 345 categories distributed among six domains. Moreover, based on the Semantic Shift Benchmark (SSB) Vaze et al. (2021) (including CUB Welinder et al. (2010), Stanford Cars Krause et al. (2013b), and FGVC-Aircraft Maji et al. (2013)), we construct a new corrupted dataset called SSB-C (i.e., CUB-C, Scars-C, and FGVC-C) following Hendrycks & Dietterich (2019). We exclude unrealistic corruptions and corruptions that may lead to domain leakage to ensure that the model does not see any of the domains in SSB-C during training (see Appendix A for details). Overall, we introduce 9 types of corruption and 5 levels of corruption severity for each type, resulting in a dataset 45$\times$ larger than SSB.
@@
For the semantics axis, on both DomainNet and SSB-C, following Vaze et al. (2022), we sample a subset of all classes as the old classes and use 50% of the images from these labelled classes to construct $\mathcal{D}_{\Omega^{a}}^{l}$. The remaining images with both old classes and new classes are treated as the unlabelled data $\mathcal{D}_{\Omega^{a}}^{u}$. For the domain axis, on DomainNet, we select images from the ‘real’ domain as $\mathcal{D}_{\Omega^{a}}$ and pick one of the remaining domains as $\mathcal{D}_{\Omega^{b}}$ in turn (or include all the remaining domains as $\mathcal{D}_{\Omega^{b}}$). While on SSB-C, we use each dataset in SSB as $\mathcal{D}_{\Omega^{a}}$ and its corresponding corrupted dataset in SSB-C as $\mathcal{D}_{\Omega^{b}}$. Statistics of the datasets are shown in Table 1.
@@
Evaluation protocol. For DomainNet, $\Omega^{a}=\left\{\omega_{1}\right\}$ and $\Omega^{b}=\left\{\omega_{2}\right\}$, where $\omega_{i}$ stands for different domains. We also experiment with the case where $\Omega^{b}=\left\{\omega_{2},\cdots,\omega_{6}\right\}$. We train the models on $\mathcal{D}_{\Omega^{a}}$ (i.e., $\mathcal{D}_{\Omega^{a}}^{l}\cup\mathcal{D}_{\Omega^{a}}^{u}$) and $\mathcal{D}_{\Omega^{b}}^{u}$ of all classes without annotations. For SSB-C, $\Omega^{a}=\left\{\omega_{1}\right\}$ and $\Omega^{b}=\left\{\omega_{2},\cdots,\omega_{10}\right\}$ since we have nine types of corruptions. During evaluation, we compare the ground-truth labels $y_{i}$ with the predicted labels $\hat{y}_{i}$ and measure the clustering accuracy as $ACC=\frac{1}{|D^{u}|}\sum_{i=1}^{|D^{u}|}\mathds{1}(y_{i}=\phi(\hat{y}_{i}))$, where $\phi$ is the optimal permutation that matches the predicted cluster assignments to the ground-truth labels. We report the ACC values for ‘All’ classes (i.e., instances from $\mathcal{Y}$), the ‘Old’ classes subset (i.e., instances from $\mathcal{Y}^{l}$), and ‘New’ classes subset (i.e., instances from $\mathcal{Y}^{u}$) for $\mathcal{D}_{\Omega^{a}}^{u}$ and $\mathcal{D}_{\Omega^{b}}^{u}$ separately.
@@
Implementation details. Following the common practice in GCD, we use the DINO Caron et al. (2021) pre-trained ViT-B/16 as the feature backbone and the number of categories is known as in Wen et al. (2023) for all methods for fair comparison. When the category number is unknown, one can employ existing methods (e.g., Han et al. (2019); Vaze et al. (2022); Hao et al. (2023); Zhao et al. (2023)) to estimate it and substitute it into the category discovery methods (see Table 24). The 768-dimensional embedding vector corresponding to the CLS token is used as the image feature. For the feature backbone, we only fine-tune the last Transformer layer. We train each dataset for $T=200$ epochs using a batch size of 256. We follow the protocol in  Vaze et al. (2022); Wen et al. (2023) to select the optimal hyperparameters for our method and all baselines, based on the ‘All’ accuracy on the validation split of $\mathcal{D}_{\omega_{1}}$. The initial learning rate for our approach is 0.1 for CUB and 0.05 for other datasets, and the rate is decayed using a cosine schedule. $t^{\prime}$ is set to the $80$-th epoch. $r_{0}$ is assigned as $|\mathcal{D}^{l}|/|\hat{\Omega}^{b}|$ for DomainNet and 0 for SSB-C. $r^{\prime}$ is set to 1 for DomainNet and 0.05 for SSB-C. $\epsilon$ is set to 0.1. Following Vaze et al. (2022); Wen et al. (2023), we set $\lambda=0.35$. See Appendices M and N for choices of hyperparameters for HiLo components and learning rates for all methods.
@@ 4.2 Main comparison
We compare our method with ORCA Cao et al. (2021), GCD Vaze et al. (2022) and SimGCD Wen et al. (2023) in generalized category discovery, along with two strong baselines RankStats+ Han et al. (2021) and UNO+ Fini et al. (2021) adapted from novel category discovery, on DomainNet (Table 2) and SSB-C (Table 3), respectively. Additionally, we provide results by incorporating UDA techniques in Section 4.3 and the strong CLIP model in Appendix G.
In Table 2, we present results on DomainNet considering one domain as $\Omega^{b}$ each time. Our method consistently outperforms other methods for ‘All’ classes (even better for ‘New’ classes) in both domain $\Omega^{a}$ and $\Omega^{b}$ by a large margin. For example, for the ‘Real’ and ‘Painting’ pair, it outperforms the GCD SoTA method, SimGCD, by nearly 5% and 19% in proportional terms, which is remarkable considering the gap between different methods. RankStat+ performs well on ‘Old’ categories in the unseen domain $\Omega^{b}$. In Appendix B, we present results on DomainNet considering all domains except ‘Real’ as $\Omega^{b}$. It can be seen, in such a challenging mixed domain scenario, our method still substantially outperforms other methods. A breakdown evaluation of each domain shift for both datasets can be found in Appendices H and I.
@@
Table 3: Evaluation on SSB-C datasets. We report results of baselines in the seen domain (i.e., Original) and the overall performance of different corruptions (i.e., Corrupted). On ‘Corrupted’, our model provides between 20% and 80% relative gains over SimGCD Wen et al. (2023).
@@
Comparing the results with the single domain results in Vaze et al. (2022); Wen et al. (2023), we find that including corrupted data during training impairs the performance on the original domain. SSB-C is 45$\times$ larger than SSB, posing a significant challenge and resulting in unsatisfactory performance for existing methods. However, our method, HiLo, continues to demonstrate promising results, further validating its effectiveness.
@@
Reference SimGCD Wen et al. (2023) 61.3 77.8 52.9 34.5 35.6 33.5
(1) SimGCD + PatchMix in Zhu et al. (2023) 62.5 76.3 54.2 34.8 36.0 33.8
@@
(5) $\bm{z}_{d},\bm{z}_{s}$ from deep features only 28.2 40.3 22.7 13.6 20.0 11.0
(6) $\bm{z}_{d},\bm{z}_{s}$ from shallow features only 10.1 18.1 6.4 5.7 9.2 5.7
(7) Self-dist. for domain head 63.2 76.8 56.1 40.2 40.5 39.8
@@ 4.3 Analysis
Effectiveness of different components. We validate the effectiveness of different components and design choices for our method in Table 4. As our method is built upon SimGCD, the effectiveness of each component can be observed by comparing its performance with that of SimGCD. We combine SimGCD with the original PatchMix in Zhu et al. (2023) (row 4) as a strong baseline for our task since these are SoTA methods for GCD and UDA respectively. Rows 4-4 indicate our main conceptual methodological contributions. As can be seen, simply combining SimGCD with the original PatchMix developed for UDA leads to a relatively small influence on the results. The original PatchMix focuses mainly on bridging the domain gap of labelled classes through a semi-supervised loss, which limits its capability on the unseen classes from new domains. After subtly adapting PatchMix into contrastive learning for GCD (row 4), the unlabelled data containing both domain shifts and semantic shifts can be properly utilized for training, leading to an obvious performance boost on $\Omega^{b}$. Furthermore, when we disentangle semantic features from domain features (row 4), the model significantly improves performance on both $\Omega^{a}$ and $\Omega^{b}$, demonstrating dissociation of spurious correlations.  Appendix E also shows the efficacy of MI regularization in two distinct scenarios. Curriculum sampling further enhances performance on $\Omega^{b}$ (row 4).
Incorporating various techniques for GCD with domain shifts. We study the effectiveness of incorporating the SoTA UDA techniques (MCC and NWD) into the baseline methods. Differently to our task, domain adaptation does not consider discovering categories in the unlabelled images from the unseen domain. Results are shown in Table 5. By comparing the results with those of ‘Real + Painting’ in Table 2, we can see that the results for each method are marginally improved by introducing these techniques. This reveals that simply adopting the UDA techniques to the GCD methods is not sufficient to handle the challenging problem of GCD with domain shifts. Moreover, HiLo again notably outperforms all other methods after introducing these UDA techniques, despite the gain by these techniques being relatively marginal, further demonstrating the effectiveness and significance of our HiLo design. We also extend our analysis by incorporating various UDA techniques (e.g., EFDM Zhang et al. (2022), SFA Li et al. (2021), UniOT Chang et al. (2022)), data augmentation methods (e.g., Mixstyle, Mixup, Cutmix) and curriculum learning (e.g., CL Bengio et al. (2009), SPL Kumar et al. (2010)) into baseline models. Our findings reveal that while some UDA techniques and data augmentations offer improvements, they fall short of addressing the full complexity of GCD with domain shifts. Specifically, EFDM improves SimGCD’s performance only on $\Omega^{a}$, likely due to its reliance on explicit source-target domain alignment, which is not available in our task formulation. The performance of different augmentation methods has a clear drop when compared with our proposed PatchMix for CL (Table 4 row 4). Notably, HiLo consistently outperforms all tested UDA baselines and data augmentation techniques. These results underscore the necessity of our tailored approach for the challenging task of GCD with domain shifts, demonstrating that simply adopting existing UDA or data augmentation methods is insufficient to address this complex problem.
Real Painting
Methods All Old New All Old New
RankStats+ +MCC+NWD 37.3 62.1 23.4 31.0 51.2 9.2
UNO+ 46.9 72.4 32.8 32.1 47.6 17.7
ORCA 33.4 50.1 26.7 30.0 41.1 9.1
GCD 50.6 54.0 48.4 34.0 43.1 22.7
SimGCD 63.1 77.1 56.9 35.7 39.0 32.4
HiLo (Ours) 65.0 77.8 58.0 42.5 43.1 42.0
Table 6: Influence of different UDA techniques (e.g., Mixstyle, EFDM), data augmentations (e.g., Mixup, Cutmix), and curriculum learning (e.g., CL, SPL). We select the ‘Real’ and ‘Painting’ domains from DomainNet to train the DINO model with the SoTA GCD method SimGCD as the baseline and compare with one baseline method UniOT from universal domain adaptation.
Real Painting
Methods All Old New All Old New
Reference SimGCD 61.3 77.8 52.9 34.5 35.6 33.5
SimGCD + Mixstyle 62.3 76.8 54.0 35.0 36.1 34.0
SimGCD + EFDM 62.6 76.0 54.7 34.1 34.8 33.8
SimGCD + Mixup 62.7 76.5 54.3 34.9 37.2 32.5
SimGCD + Cutmix 62.5 76.3 54.1 33.2 36.0 31.6
Reference HiLo 64.4 77.6 57.5 42.1 42.9 41.3
HiLo + Mixup - PatchMix 63.7 77.1 56.8 39.8 39.9 38.7
HiLo + Cutmix - PatchMix 62.9 76.8 56.0 37.4 38.0 36.7
HiLo + CL - curriculum sampling 62.0 75.9 53.2 34.7 35.8 33.8
HiLo + SPL - curriculum sampling 62.8 76.5 54.5 35.0 36.1 34.0
Reference HiLo 64.4 77.6 57.5 42.1 42.9 41.3
SFA 60.1 73.4 52.8 34.9 38.0 32.6
UniOT 63.3 77.4 57.4 35.3 38.7 32.0
@@ Acknowledgements
This work is supported by National Natural Science Foundation of China (Grant No. 62306251), Hong Kong Research Grant Council - Early Career Scheme (Grant No. 27208022), and HKU Seed Fund for Basic Research.
@@ Appendix A SSB-C benchmarks
As demonstrated in Section 4.1 in the main paper, we construct the SSB-C benchmark to evaluate the robustness of algorithms to diverse corruptions applied to validation images of the original SSB benchmark (including CUB Welinder et al. (2010), Stanford Cars Krause et al. (2013a) and FGVC-Aircraft Maji et al. (2013)), adopting the corruptions following Hendrycks & Dietterich (2019). We introduce 9 types of corruption (see Figure 4) in total. Each type of corruption has 5 severity levels. Therefore, SSB-C is 45$\times$ larger than the original SSB.
@@
Here is the list of the 9 types of corruption we applied following Hendrycks & Dietterich (2019):
@@
Though synthetic, SSB-C incorporates extra challenges and unique values over existing datasets like DomainNet Peng et al. (2019a). Particularly, SSB-C includes (1) fine-grained recognition challenges under domain shifts and (2) more types of domain shifts that are not covered in DomainNet.
@@ Appendix D Theoretical analysis
Recall that $\mathcal{D}$ is an open-world dataset consisting of a labelled set $\mathcal{D}^{l}=\{(\bm{x}_{i},y_{i})\}_{i=1}^{N_{l}}\subset\mathcal{X}^{l}\times\mathcal{Y}^{l}$ and an unlabelled set $\mathcal{D}^{u}=\{\bm{x}_{i}\}_{i=1}^{N_{u}}\subset\mathcal{X}^{u}$. We also define the mapping function $g$ parametrized by a deep neural network as one of the hypotheses from $\mathbb{G}$.
@@
When the source domain does not adequately cover the target domain, the target risk of a learned hypothesis cannot be consistently estimated without additional assumptions. Nonetheless, an upper bound on the target risk can be estimated and then minimized. Ben-David et al. (2006) introduce the $\mathcal{A}$-distance (also known as $\mathcal{H}$-divergence) to assess the worst-case loss when extrapolating between domains for hypothesis classes. The $\mathcal{A}$-distance between any two distributions $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ is defined as